Please help. Also, I really don't get this part. W
Please help. Also, I really don't get this part. W
Please help. Also, I really don't get this part. Why bond therapy of H is 502kJ in H2O?
Answer #1

1) Given reaction O3 + NO ------------> NO2 + O2

O3 contains O=O and O-O

NO contains N=O

NO2 contains N=O and N-O

O2 contains O=O.

Then,

ΔH = Sum of bond energy of reactants - Sum of bond energy of products

=  bond energy of O3 +  bond energy NO - {bond energy of NO2 + bond energy of O2 }

= bond energy of (O=O + O-O) +  bond energy of N=O - {bond energy of (N=O + N-O) + bond energy of O=O }

= 498 kJ/mol + 142 kJ/mol + 590 kJ/mol - {590 kJ/mol + 222 kJ/mol + 498 kJ/mol}

= -80 kJ/mol

ΔH =  -80 kJ/mol

2) Given reaction H2O --------> H + OH   ΔHo = 502 kJ

This is not H-H bond energy.

This is H-OH bond enthalpy.